OpenStudy (anonymous):

how many gallons of elastometric paint are requires to paint the exterior surface area of an open cylindrical concrete tank of height 3.6m and base outside diameter of 2.0m if one gallon of paint covers 10m^2

OpenStudy (anonymous):

Finding out the area of the tank \[A = 2*\pi*r*h + \pi*r^2\]\[A = 2*\pi*1*3.6 + \pi*1^2\]\[A=8.2*\pi\] Number of gallons \[n = \frac{ 8.2*\pi }{ 10 }\]\[n = 2.57 gallons\] or (if the answer must be in integer) \[n = 3 gallons\]

OpenStudy (anonymous):

thankyou :)